# Efficient Training Spiking Neural Networks with Parallel Spiking Unit

### v2

@@ III-A Vanilla LIF Neuron
Here, $V[t]$ is the membrane potential of the neuron, which adjusts itself with the synaptic current $I[t]$. When the neuron’s membrane potential exceeds the threshold $V_{th}$, it fires a spike. Here, $\Theta$ is the Heaviside function. There are two forms of neuron reset: hard reset and soft reset[36], differing in whether the membrane potential is cleared after spike firing to further receive subsequent information.
@@ III-B Parallel Spiking Unit
The reset process in the LIF model diminishes the membrane potential’s strength, essentially causing some of the historically integrated information to be forgotten. This reset mechanism depends on the neuron’s spike activity at each time step. Focusing on the soft reset scenario, the membrane potential of a neuron that fires a spike at time $t$ will be adjusted by $-(1-\frac{1}{\tau})V_{th}S_{i}[t]$ at time $t+1$. The leakage coefficient $\tau$ ensures that this loss of information is further modified to $-(1-\frac{1}{\tau})^{2}V_{th}S_{i}[t]$ at time $t+2$, and this pattern continues over time. Consequently, the information loss $[R[1],R[2],\ldots,R[T]]$ induced by the reset process at each time step can be computed in parallel:
@@
In contrast to SPSN, the PSU model omits the Bernoulli process and employs a deterministic approach to spiking, which does not impede the computation of potential events in $I=WX$. Spike estimation relies exclusively on the membrane potential $S^{\prime}=AI$, whereas the actual spike discharge during the reset phase is represented by $AI-BS$, with $S$ denoting the actual spike activity. Given that the interaction between $BS$ involves matrix multiplication, the spike discharge rate at subsequent time steps tends to be elevated. Nevertheless, utilizing $S^{\prime}$ in the Reset computation reduces the firing rate compared to the traditional LIF model, offering benefits for sparse computation within SNNs. Besides setting the hyperparameter $\tau$ similar to the LIF model, the PSU does not involve any additional hyperparameters. Its computation process exhibits higher sparsity, and the employment of parallel computation markedly enhances training and inference efficiency. Theoretically, compared to the traditional LIF neuron model, the PSU demonstrates a time complexity of $\mathcal{O}(1)$ in the inference process.
@@ IV-A Experimental Setup
CIFAR10, a well-known dataset in machine learning, is used for benchmarking image recognition algorithms. It contains 60,000 32x32 color images in ten classes. Sequential CIFAR10/100 datasets present images in sequences to train models on temporally related patterns. DVS datasets, like DVSCIFAR10, are created with neuromorphic vision sensors, capturing scene changes asynchronously and sparsely, suitable for neuromorphic visual data processing. The SHD speech dataset is tailored for spiking neural networks, featuring spoken digits in English and German, making it bilingual.
@@ IV-C Neural Activity Analysis
Beyond the speed advantage in simulation, we expect that the computational processes in spiking neurons will demonstrate greater sparsity, potentially further diminishing the energy consumption required for computations in the SNN upon deployment. Our analysis of the spike issuance rate within the initial eight layers of VGGSNN during the classification of DVSCIFAR10, as illustrated in Figure 5(a), supports this anticipation. The data suggests that the PSU architecture can conduct computations with reduced spike activity. In contrast, the IPSU model, which utilizes learnable parameters to adjust the membrane potential $V$, shows increased spike activity following the learning phase. On the other hand, RPSUs, which integrate learnable parameters within the Sigmoid function, exhibit more consistent learning outcomes and enhanced handling of dynamic information interactions while maintaining a lower spike activity level than that observed in the LIF model. Moreover, despite its rapid simulation capabilities, the SPSN model demonstrates instability in neural activity and proves inefficient training due to the lack of a reset mechanism. Similarly, the PLIF model, which actively and adaptively adjusts leakage coefficients in response to incoming data, also does not achieve the level of sparsity seen in the PSU model. In light of the observation that PSU and RPSU can maintain comparable or superior performance with less frequent spiking activity, we delved deeper into this phenomenon. We apply these models to identical data samples and visualized neuron activities in the first and third layers, as depicted in Fig. 5(b). The findings highlight that PSU and its derivatives are adept at minimizing non-essential features’ representation while preserving pivotal elements. This starkly contrasts the IPSU model, which diverges from the LIF model in its learning approach, leading to a unique pattern in spiking activity.
@@ V Conclusion
The acceleration of spiking units is a topic of interest but directly determines the efficiency of SNN applications. This paper presents a significant advancement in SNNs by developing Parallel Spiking Units (PSUs) and their variants, IPSU and RPSU. By innovatively decoupling the leaky integrate-and-fire mechanism and introducing a probabilistic reset process, this approach enables parallel computation of neuron membrane potentials, significantly enhancing computational efficiency and speed. Our findings demonstrate that PSU, IPSU, and RPSU outperform traditional LIF neurons and existing spiking neuron models in various datasets, offering sparser spiking activity and faster processing. Importantly, this study addresses the critical limitation of SNNs in utilizing parallel computing resources like GPUs, paving the way for low-power, high-performance AI systems. This breakthrough marks a substantial contribution to neuromorphic computing, opening new avenues for applying and developing advanced SNNs in real-world scenarios.

### v3

# Parallel Spiking Unit for Efficient Training of Spiking Neural Networks

@@ III-A Vanilla LIF Neuron
Here, $V[t]$ is the membrane potential of the neuron, which adjusts itself with the synaptic current $I[t]$. When the neuron’s membrane potential exceeds the threshold $V_{th}$, it fires a spike. Here, $\Theta$ is the Heaviside function. Two forms of reset are considered: hard reset and soft reset[36], differing in whether the membrane potential is cleared after spike firing to further receive subsequent information.
@@ III-B Parallel Spiking Unit
The reset process in the LIF model diminishes the membrane potential’s strength, essentially causing some of the historically integrated information to be forgotten. This reset mechanism depends on the neuron’s spike activity at each time step. Focusing on the soft reset scenario, the membrane potential of a neuron that fires a spike at time $t$ will be adjusted by $-(1-\frac{1}{\tau})V_{th}S[t]$ at time $t+1$. The leakage coefficient $\tau$ ensures that this loss of information is further modified to $-(1-\frac{1}{\tau})^{2}V_{th}S[t]$ at time $t+2$, and this pattern continues over time. Consequently, the information loss $[R[1],R[2],\ldots,R[T]]$ induced by the reset process at each time step can be computed in parallel:
@@
In contrast to SPSN, the PSU model omits the Bernoulli process and employs a deterministic approach to spiking, which does not impede the computation of potential events in $I=WX$. Spike estimation relies exclusively on the membrane potential $S^{\prime}=AI$, whereas the actual spike discharge during the reset phase is represented by $AI-BS$, with $S$ denoting the actual spike activity. Given that the interaction between $BS$ involves matrix multiplication, the spike discharge rate at subsequent time steps tends to be elevated. Nevertheless, utilizing $S^{\prime}$ in the Reset computation reduces the firing rate compared to the traditional LIF model, offering benefits for sparse computation within SNNs. Besides setting the hyperparameter $\tau$ similar to the LIF model, the PSU does not involve any additional hyperparameters. Its computation process exhibits higher sparsity, and the employment of parallel computation markedly enhances training and inference efficiency.
@@ IV-C Neural Activity Analysis
Beyond the speed advantage in simulation, we expect that the computational processes in spiking neurons will demonstrate greater sparsity, potentially further diminishing the energy consumption required for computations in the SNN upon deployment. Our analysis of the spike issuance rate within the initial eight layers of VGGSNN during the classification of DVSCIFAR10, as illustrated in Fig. 5(a), supports this anticipation. The data suggests that the PSU architecture can conduct computations with reduced spike activity. In contrast, the IPSU model, which utilizes learnable parameters to adjust the membrane potential $V$, shows increased spike activity following the learning phase. On the other hand, RPSUs, which integrate learnable parameters within the Sigmoid function, exhibit more consistent learning outcomes and enhanced handling of dynamic information interactions while maintaining a lower spike activity level than that observed in the LIF model. Moreover, despite its rapid simulation capabilities, the SPSN model demonstrates instability in neural activity and proves inefficient training due to the lack of a reset mechanism. Similarly, the PLIF model, which actively and adaptively adjusts leakage coefficients in response to incoming data, also does not achieve the level of sparsity seen in the PSU model. In light of the observation that PSU and RPSU can maintain comparable or superior performance with less frequent spiking activity, we delved deeper into this phenomenon. We apply these models to identical data samples and visualized neuron activities in the first and third layers, as depicted in Fig. 5(b). The findings highlight that PSU and its derivatives are adept at minimizing non-essential features’ representation while preserving pivotal elements. This starkly contrasts the IPSU model, which diverges from the LIF model in its learning approach, leading to a unique pattern in spiking activity.
@@ V Conclusion
The acceleration of spiking units is a topic of interest but directly determines the efficiency of SNN applications. This paper presents a significant advancement in SNNs by developing Parallel Spiking Units and their variants, IPSU and RPSU. By innovatively decoupling the leaky integrate-and-fire mechanism and introducing a probabilistic reset process, this approach enables parallel computation of neuron membrane potentials, significantly enhancing computational efficiency and speed. Our findings demonstrate that PSU outperforms traditional LIF neurons and existing spiking neuron models in various datasets, offering sparser spiking activity and faster processing. Importantly, this study addresses the critical limitation of SNNs in utilizing parallel computing resources like GPUs, paving the way for low-power AI systems.
Acknowledgment
This research was financially supported by Postdoctoral Fellowship Program of CPSF (Grant No. GZC20232994) and a funding from Institute of Automation, Chinese Academy of Sciences (Grant No. E411230101).